OpenStudy (anonymous):

Can someone please help! I GIVE MEDALS!! The equation below can be used to find the length of a foot or forearm when you know one or the other. (length of the foot) = 0.860 • (length of the forearm) + 3.302 If you let y = length of the foot and x = length of the forearm, this equation can be simplified to y = 0.860x + 3.302. Using this equation, how long would the foot of a person be if his forearm was 17 inches long?

OpenStudy (anonymous):

Using this equation, how long would the foot of a person be if his forearm was 17 inches long? What is the rate of change of the equation from Part A? Compare the equation from Part A to your data. Are they the same? Which has a greater rate of change? Why do you think the values are different? Is the relation in your data a function? Why or why not? Could the equation in Part A represent a function? Why or why not? Explain your answer.

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

given eq y =0.860x +3.302 this eq is given in the slope form of a line y = mx+c m = rate of change put x = 17 to find y

OpenStudy (anonymous):

y=17.922

OpenStudy (anonymous):

I think?

OpenStudy (anonymous):

Part A = y = 0.860x + 3.302 y = 0.860 * 17 + 3.302 = 19.02

OpenStudy (anonymous):

Can you help me on the second part too? I really need help.

OpenStudy (anonymous):

rate of change = m = 0.860

OpenStudy (anonymous):

what does the 'm' mean?

OpenStudy (anonymous):

slope

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

what about the rest of it : Compare the equation from Part A to your data. Are they the same? Which has a greater rate of change? Why do you think the values are different? Is the relation in your data a function? Why or why not? Could the equation in Part A represent a function? Why or why not? Explain your answer.

OpenStudy (anonymous):

I need help in that. Let me ask @StudyGurl14 @DanJS

OpenStudy (anonymous):

okay :)

OpenStudy (danjs):

hi

OpenStudy (anonymous):

Hi @DanJS please help

OpenStudy (anonymous):

Hello, can you please help :)

OpenStudy (danjs):

ok which question

OpenStudy (anonymous):

Compare the equation from Part A to your data. Are they the same? Which has a greater rate of change? Why do you think the values are different? Is the relation in your data a function? Why or why not? Could the equation in Part A represent a function? Why or why not? Explain your answer.

OpenStudy (danjs):

where is your data? is there a table that goes with this?

OpenStudy (anonymous):

this is what I have done: Compare rates of change: (Length of the foot) = 0.860 • (length of the forearm) + 3.302 If you let y = length of the foot and x = length of the forearm, this equation can be simplified to y = 0.860x + 3.302. Using this equation, how long would the foot of a person be if his forearm was 17 inches long? The foot would be 19.02 inches long. y = 0.860 * 17 + 3.302 = 19.02 inches long. rate of change = m = 0.860

OpenStudy (danjs):

I think there is some other data that is missing, a table or something, so you can use points from that to calculate a slope and compare that to the slope of the given equation

OpenStudy (anonymous):

oh wait!

OpenStudy (anonymous):

Input (Forearm) Output (Foot) (Me) 10 inches 10 inches ~ (My mom) 9 inches 10 inches ~ (My brother) 10 inches 11 inches (Baby brother) 4 inches 3 inches (Dad) 12inches 12 inches

OpenStudy (danjs):

ok, so you need to use the equation y-y1 = m(x-x1) where (x1,y1) are points in the table

OpenStudy (anonymous):

Rate of change- Change in forearm / change in feet Change in forearm- 10 – 9 = 1 Change in feet- 11 – 10 = 1 Rate of change- 1/1 = 1 So the rate of change would be on because it is increasing by 1.

OpenStudy (anonymous):

That equation is confusing!

OpenStudy (danjs):

For example, My mom (x,y) = (arm , foot)

OpenStudy (danjs):

so my mom(x,y) = (9,10) need to find the rate of change and compare to the original equation rate of change

OpenStudy (anonymous):

rate of change for first one - 1( is that right?)